it takes a computer 30 mins to process and print payroll checks, when a second computer is used and the two computers work together the checks can be processed and printed in 21 minutes. how long would it take the second computer by itself to process and print the payroll checks?
Found 2 solutions by jorel555, Boreal:
Answer by jorel555(1290)

Let n be the number of minutes the second computer takes. Then:
1/30+1/n=1/21
14n+420=20n
6n=420
n=70!!!!!!!!!!!!!!!!!!!!

Answer by Boreal(15235)

in one minute computer A does 1/30 the work.
in one minute computer B does 1/x the work
(1/30)+(1/x)=1/21 Both together do 1/21 in one minute
multiply everything by 630x
21x+630=30x
9x=630
x=70 minutes